### v1

# Learning phase transitions by siamese neural network

Abstract
The wide application of machine learning (ML) techniques in statistics physics has presented new avenues for research in this field. In this paper, we introduce a semi-supervised learning method based on Siamese Neural Networks (SNN), trying to explore the potential of neural network (NN) in the study of critical behaviors beyond the approaches of supervised and unsupervised learning. By focusing on the (1+1) dimensional bond directed percolation (DP) model of nonequilibrium phase transition, we use the SNN to predict the critical values and critical exponents of the system. Different from traditional ML methods, the input of SNN is a set of configuration data pairs and the output prediction is similarity, which prompts to find an anchor point of data for pair comparison during the test. In our study, during test we set different bond probability $p$ as anchors, and discuss the impact of the configurations at this anchors on predictions. More, we use an iterative method to find the optimal training interval to make the algorithm more efficient, and the prediction results are comparable to other ML methods.
Introduction
With the rapid progress of the computer field in the past decade, machine learning (ML) has been explosively applied in various fields of science [1, 2]. By the great power of feature capture and classification, ML has brought new inspiration to the study of statistical physics and proven to be an effective tool for identifying and classifying different phases [3, 4, 5, 6, 7]. Through training process, ML algorithms can learn the underlying patterns and laws of phase transition from the input configurations of a given system, enabling them to make predictions in test data. Additionally, ML algorithms can automatically extract and select the most relevant features, which are valuable for reducing the complexity of high-dimensional phase transition data.
The mainly types of ML method applied to study phase transitions are supervised learning [5, 8, 9, 10, 11], unsupervised learning [6, 12, 13, 14, 15, 16, 11], and semi-supervised learning  [17, 18, 19, 20, 21]. For supervised learning, the input data for training needs to be labeled, which commonly used to identify or classify the phases of matter. For unsupervised learning, the labeling of the order parameter is not required, such as the methods of principal component analysis (PCA)[22, 12, 16], t-distributed stochastic neighbor embedding (T-SNE) [23] or nonlinear autoencoder (AE) [23, 11, 12, 16], for clustering and dimensionality reduction. Recently, semi-supervised learning, especially transfer learning (TL) also has been proved as a good choice to detect critical points and calculate critical exponents [20], even determining the type of phase transition [21]. With the input of mixing both labeled and unlabeled data, methods of TL can transform unlabeled data in target domain into labeled data in source domain by training, which only use a limited labeled set for the purposes of prediction. Although the methods of ML are dazzling, all of them are not fully perfect. For example, supervised learning must time-consumingly label the data; unsupervised learning is hard to fully identify the dynamic features with limited data; semi-supervised learning is very sensitive to the input, which always need to be preprocessed. Therefore, it is still imperative to find a more complete and efficient method of ML to study phase transitions.
Siamese neural network (SNN) [24, 25, 26], also called twin NN, is a specialized architecture usually consisting of two parallel identical neural networks (NN) that share their weight for the purposes of assessing or comparing the similarity between two input objects. SNN is not simply classified into supervised learning, unsupervised learning or semi-supervised learning, but a flexible structure that fill them in. It is widely used in computer vision for face recognition [27], target tracking [28] and image retrieval [29], and also shows its power in natural language processing (NLP) for text matching tasks [30]. In the study of phase transition, Patel $etal$ [31] introduced a SNN technique of unsupervised learning to identify phase boundaries in Ising-type systems and Rydberg atomic arrays, showing its ability to learn about multiple phases without knowing about their existence. As the preliminary attempt, SNN is proved as a valuable tool for discovering new and unfamiliar phases of matter, and worth for further application and exploration.
Inspired by this, we apply a new SNN algorithm based on semi-supervised learning to study the critical behaviors of phase transition. In this method, we divide input configurations to two separate sets for training. By choosing an anchor for configurations, the SNN would determine whether this samples belong to the same set through training, which gives the power to judge phases and find critical point. For test, we focus on the model of (1+1)-dimensional bond directed percolation (DP) [32, 33, 34], belonging to the most important universality class of nonequilibrium phase transitions, the DP universality class [35, 36, 37]. We not only predict the critical point of the DP model through the SNN, but also calculate the critical exponent by the data collapse method [38, 39]. Further, we discuss the impact of anchor point selection on results, and find the optimal interval of training set to achieve more accurate predictions. By doing so, we hope to enhance the effectiveness and reliability of the SNN as a new tool in the study of phase transitions.
The remaining content of this paper is structured in the following manner. In Section 1, we have introduced the directed percolation model. Section 2 gives the methods of SNN and the data set for study. Section 3 shows the estimate critical probabilies of SNN and also the prediction of spatial correlation exponent $\nu_{\perp}$. In Section 4, we discuss the case of chosen archor in critical region, and the impact of the training set on the results. Section 5 is a summary of this work.
1 Model
1.1 The directed percolation model
As the most prominent universality class of absorbing phase transition, DP describes percolation of one direction along that the system falls into inactive state where the time evolution stops. It is utilized to investigate various phenomena including the propagation of epidemics [40], forest fires [41], and traffic flow  [42], etc. In past decades, a various of methods has been applied to study DP models, such as mean-field theory [43], renormalization groups [44], field theory approach [45] and numerical simulations [46] and so on. Some ML approaches have also been tried recently and shown their own strengths [20, 47]. Since the models of DP universality class include unitary or binary random reaction processes, which may contain diffusion or non-diffusion motions, only the critical exponents are common and the position of the phase transition may vary. In our work, we use a SNN based algorithm to study the $(1+1)$-dimensional bond DP, one of the simplest cases of DP universality class.
Fig. 1 illustrates the configurations of (1+1)-dimensional DP on the square lattice of size $L=500$ with the evolution time step $t=500$. Starting from a fully occupied lattice, a bond is formed at the time step with probability $p$ from an existing bond, which can be interpreted as a reaction-diffusion mechanism of interacting particles with the active particle $A$ and the empty site $\varnothing$:
$\left\{\begin{array}[]{rcl}\mbox{self-destruction:}&A\longrightarrow\varnothing,\\ \mbox{diffusion:}&\varnothing+A\longrightarrow A+\varnothing,\\ \mbox{offspring production:}&A\longrightarrow 2A,\\ \mbox{coagulation:}&2A\longrightarrow A.\\ \end{array}\right.$ (1)
(a) $p<p_{c}$
(b) $p=p_{c}$
(c) $p>p_{c}$
Figure 1: Configurations of the (1+1) dimensional bond DP corresponding to three different bond probabilities $p=0.5,0.6447$ and $0.8$ respectively.
By the arrows, this system evolves with the following rules:
$s_{i}(t+1)=\left\{\begin{array}[]{rcl}1&&{if\quad s_{i-1}(t)=1\quad\text{and}\quad z_{i}^{-}<p,}\\ 1&&{if\quad s_{i+1}(t)=1\quad\text{and}\quad z_{i}^{+}<p,}\\ 0&&\text{otherwise},\end{array}\right.$ (2)
where $s_{i}(t)$ is the state of site $i$ at time $t$ and $p$ is the bond probability of a connection between two adjacent sites. $z_{i}^{\pm}\in(0,1)$ is a random number taken from an uniform distribution. From the definitions above, the configuration of the system can be determined by the states of all sites.
The models of DP universality class also has their own independent critical exponents. The order parameter of the (1+1)-dimensional bond DP close to the critical point $p_{c}$ can be expressed by the steady-state particle density $\rho_{a}$ according to the critical exponent $\beta$ as [35],
$\rho_{a}(p)\widetilde{\propto}(p-p_{c})^{\beta}.$ (3)
In this steady state, the system also obeys the following power-law expressions of spatial and temporal correlation length:
$\xi_{\perp}\sim\mid p-p_{c}|^{-\nu_{\perp}},\quad\xi_{\parallel}\sim\mid p-p_{c}|^{-\nu_{\parallel}},$ (4)
where $\nu_{\perp}$ and $\nu_{\parallel}$ are the spatial and temporal correlation exponent.
In this paper, we use the the full configurations of the (1+1)-dimensional DP shown in Fig. 1 as input to the SNN algorithm. To generate the configurations, we use periodic boundary conditions, and originally starts with fully occupied lattice (explained in our previous study [20]). The characteristic temporal length $t_{c}\sim L^{z/d}$ is set by the dynamical exponent $z=1.580(1)$ and spatial dimension $d=1$, to ensure the stable results [35].
2 Method
2.1 The Siamese Neural Network (SNN) method
The SNN is a type of deep learning model commonly used for tasks involving metric learning and similarity comparison [24]. In various applications, SNNs are frequently employed to analyze pairs of input data, such as images, text, or voice [24]. It works by creating multiple parallel sub-networks, known as "Branches", which share the same structure and weights. By feeding each input data through the NN of these branches for feature extraction, the SNN can calculate the similarity scores between them using metric functions like Euclidean distance or cosine similarity. The significant advantage of SNN is the capability to acquire adaptable feature representations while training, which enables the NN to achieve high performance even on datasets having some degree of variability. Additionally, SNN can be trained end-to-end by back-propagation algorithms, which allows the entire NN to autonomously learn feature representations and measures of similarity.
Figure 2: The structure of SNN
The specific framework of SNN used in our work to study phase transition of the (1+1)-dimensional DP is shown in the Figure 2. We use the raw configurations of $L=8,16,32,48$ and $64$ as input, and the details of dataset selection will be given in Section 2.2. The input is a set of data pairs $(S_{i},S_{j})$, $S_{i}$ enters one branch in the framework, and $S_{j}$ enters the other. $S_{i}$ and $S_{j}$ are two different configuration samples of the same type of data. Taking the branch with $S_{i}$ as an example, the input $S_{i}$ is flattened into one-dimensional data and fed into an fully connected (FC) layer of $500$ neurons with a $ReLu$ activation function. Here we also set a Batch Normalization (BN) to avoid over-fitting and speed up the NN to optimize its parameters. The output of this FC layer is a feature representation $f_{1}$ of length $50$. Therefore, the entire calculation process of the first branch can be expressed as function $f_{1}(S_{i})$. Similarly, the second branch with $S_{j}$ has the same process, and the feature representation is recorded as $f_{2}(S_{j})$. Note that these two branches shares the weights for learning.
The shared weights enable embedded data to be in the same space from any branches. Therefore, it allows to calculate the “distance" between two the feature representations $f_{1}(S_{i})$ and $f_{2}(S_{j})$, to compare the similarity of $S_{i}$ and $S_{j}$. Here we define the “distance" by $L_{1}$ norm. Then this information of “distance" is fed into a dense layer of FC with the numbers of neurons $50$ to $1$, and this process is also attached with a $ReLu$ activation function and BN. By applying a $Sigmoid$ activation function to the output of FC, the result is converted into a probability of $[0,1]$, which is regarded as classification score $y$, for similarity.
2.2 Data sets of the SNN
For input data of the SNN, we use the full configurations of (1+1) dimension DP (time evolution step as the characteristic temporal length.). In each lattice size $L=8,16,32,48$ and $64$, we select $101$ values of $p\in[0,1]$ uniformly and generate $1000$ samples of configuration at each $p$. To minimize human intervention during training, we label the samples of configuration far away from the critical regime, i.e., the input configurations $S_{i}$ and $S_{j}$ are generated in the range of $p\in[0,0.1]\cup[0.9,1]$. The idea of labeling here is very different from the input of TL in Ref. [20]. The labels “1" or “0" indicates the input data pair $S_{i}$ and $S_{j}$ belong to the same or different phases to ensure maximum similarity or dissimilarity between them. Specifically, if $S_{i}$ and $S_{j}$ both belong to $[0,0.1]$ (or $[0.9,1]$), the input $(S_{i},S_{j})$ is labeled as “1". Or $S_{i}$ and $S_{j}$ are in the range of $[0,0.1]$ and $[0.9,1]$ respectively, the data pair is labeled as “0". For training, the loss function is defined as:
$L=\sum\limits_{i=1}^{N}y(S_{i},S_{j})\log s(S_{i},S_{j})+\big{(}1-y(S_{i},S_{j})\big{)}\log\big{(}1-s(S_{i},S_{j})\big{)},$ (5)
where $s$ is the similarity of ($S_{i}$, $S_{j}$) , and $y$ is the label, ranging from 0 to 1. The training epoch is set to $5000$, and our work in on python 3.10 and RTX 4090 Gpu partform with tensorflow 2.14.0 library.
For the process of test, a reference point, called "anchor" point, is selected to evaluate the similarity $s$ or dissimilarity $(1-s)$ of all configurations. The anchor $p_{a}$ can be selected arbitrarily within the range of $p\in[0,1]$. By taking the samples of configuration at the anchor $p_{a}$ as $S_{j}$ and the samples of the training set as $S_{i}$, in this way the similarity between the samples at the anchor and the training samples can be estimated through the SNN. And in detail the dependence of SNN results on anchor point will be discussed in Section 3.2. Then the output of SNN, $P_{0}$, can be expressed as two curves within bond probability $p\in[0,1]$: the curve $s$ of similarity and $1-s$ of dissimilarity. The intersection of these two curves or $p$ at $P_{0}=50\%$ represents the critical point of phase transition in the system.
3 Results
3.1 Test of the SNN framework
Figure 3: The output of SNN with the test anchor $p_{a}=0$ at $L=32$. The blue curve represents the similarity ($s$), while the ensemble average of the test samples is $1000$. The light blue curve represents the non-similarity ($1-s$).
According to the previous section, the entire SNN framework has been well trained by the set of samples within $p\in[0,0.1]\cup[0.9,1]$. To test the usability of SNN, firstly we select all samples of $p\in[0,0.1]$ as $S_{i}$ and only the configurations of anchor $p_{a}=0$ as $S_{j}$ at $L=32$ to form the input of test. The output $P_{0}$ is shown in the figure 3. For the similar curve $s$, $P_{0}$ is a stable value $1$ at $p\in[0,0.48]$ and 00 at $p\in[0.64,1]$, which means all configurations of anchor $p_{a}=0$ are $100\%$ having the same phase with samples of $p\in[0,0.48]$ but totally different with the samples of $p\in[0.64,1]$. In the area of $p\in(0.48,0.64)$, the similarity curve $s$ drops sharply, that is, the samples begin to be dissimilar to the ones of anchor $p_{a}=0$, and then it is completely in differ near $p=0.64$. The result of dissimilarity curve $(1-s)$ is opposite to the similarity. The evaluated critical point of SNN at $L=32$ is $p_{c}\simeq 0.6317$, which is close to the theoretical value $0.6447$ [35]. This shows that the SNN can relatively accurately predict the critical point of the (1+1) dimensional bond DP at $L=32$ by the selected anchor $p_{a}=0$
3.2 The dependence of anchor selection
(a) $L=16$
(b) $L=32$
(c) $L=48$
(d) $L=64$
Figure 4: The output of SNN with the test anchor $p_{a}=0,0.4,0.8$ and $1$, at different sizes (a) $L=16$, (b) $L=32$, (c) $L=48$ and (d) $L=64$.
Figure. 4 shows the similarity analysis of SNN with different anchors at $L=16,32,48$ and $64$. To study the dependence of anchors, we keep samples of $S_{i}$ from $p\in[0,1]$ but select anchor $p_{a}=0,0.4,0.8$ and $1$ for $S_{j}$ for comparison. Since during training process the SNN learns nothing about the samples of $p_{a}=0.4$ and $0.8$, here it also contains the idea of transfer learning, i.e., the range $p\in[0,0.1]\cup[0.9,1]$ as “source domain" and $p_{a}=0.4$ or $0.8$ as “target domain". As shown in Figure. 4, the similar curve $s$ changes with the different anchors. For $p_{a}=0$ (blue curve) and $0.4$ (brown curve), the similarity curve $s$ of all configurations is from “1" to “0" at $p\in[0,1]$, but for $p_{a}=0.8$ (red curve) and $1$ (green curve), it becomes the opposite trend as from “0" to “1". Because the configurations of $p_{a}=0$ and $0.4$ have similar phase to the ones of $p\in[0,0.1]$, but totally different at $p_{a}=0.8$ and $1$.
3.3 Extrapolated results of critical points
(a) $p_{a}=0$
(b) $p_{a}=0.4$
(c) $p_{a}=0.8$
(d) $p_{a}=1$
Figure 5: Extrapolation of the critical probability $p_{c}$ to infinite lattice size with anchor (a) $p_{a}=0$, (b) $p_{a}=0.4$, (c) $p_{a}=0.8$ and (d) $p_{a}=1$. The predictions of this $p_{c}^{\infty}$ are summarized in Table. 1.
To deeply analyze the dependence of the critical point predicted by SNN on anchor points, we extrapolate the results at $L=8,16,32,48$ and $64$ to zero on the $1/L$ scale under anchor points $p_{a}=0,0.4,0.8$ and $1$, by finite-size scale fitting (FSS) [48, 49, 50, 51]. As shown in Figure. 5, the critical values increase with the size of $L$, which is consistent with the layer of phase transition models under limited scale. The SNN results of $p_{c}$ at infinite lattice size are summarized in Table. 1. Comparing with the theoretical value $p_{c}=0.6447$ [35], it shows the SNN can relatively predict critical value $p_{c}^{\infty}$ within deviation. The closest result of SNN is $p_{c}=0.6426$ at anchor $p_{a}=0.4$, and the SNN results increase with the value of selected anchors, which also indicates the results of SNN may have anchor dependence. However, this achor dependence can be diluted by expanding the training set, which will be discussed in Subsection 4.2.
Further, the curves of $P_{0}$ in Figure. 4 can be fitted with a sigmoid function:
$p\to\frac{1}{1+e^{-\frac{(p-p_{c})}{\sigma}}}\quad\text{or}\quad p\to 1-\frac{1}{1+e^{-\frac{(p-p_{c})}{\sigma}}},$ (6)
for $1-s$ or $s$, where $\sigma$ is the scaling width. As shown in Figure. 6, by the technique of data collapse we can obtain the spatial correlation exponent $\nu_{\perp}$ via the scaling $(p-p_{c})L^{1/\nu_{\perp}}$. From $P_{0}$ at lattice sizes of $L=8,16,32,48$ and $64$, a proper value of $\nu_{\perp}$ can be found to fit the scaling. Table. 1 shows the fitted $\nu_{\perp}$ of each fixed anchors, all of which are consistent with the theoretical value $1.09$ [35]. Since $\nu_{\perp}$ has more dependence on raw configuration and also has lower precision, it is not as sensitive to changes with the anchors as $p_{c}^{\infty}$.
(a) $p_{a}=0$
(b) $p_{a}=0.4$
(c) $p_{a}=0.8$
(d) $p_{a}=1$
Figure 6: The collapse of the average output layer as a function of $(p-p_{c})L^{1/\nu_{\perp}}$, with anchor (a) $p_{a}=0$, (b) $p_{a}=0.4$, (c) $p_{a}=0.8$ and (d) $p_{a}=1$.
Table 1: The critical value $p_{c}^{\infty}$ and spatial correlation exponent $\nu_{\perp}$ estimated by SNN with archor $p_{a}=0,0.4,0.8$ and $1$.
$p_{a}=$0.0 $p_{a}=$0.4 $p_{a}=$0.8 $p_{a}=$1.0
   $p_{c}^{\infty}$ 0.6411 0.6426 0.6491 0.6508
   $\nu_{\perp}$ 1.09 1.08 1.1 1.09
4 Discussion
4.1 The anchor point $p_{a}=0.64$
Figure 7: The output of SNN with anchor $p_{a}=0.64$ at $L=8,16,32,48$ and $64$.
[…]
4.2 The optimal training set
In previous tests, we sampled configurations $S_{i}$ and $S_{j}$ in the range of $p\in[0,0.1]\cup[0.9,1]$ as the training set of SNN, to minimize human intervention. However, from Section 3 it is obviously that the expected critical region is very narrow, which makes it possible to expand the training set. Typically, a wider and more accurate label training set is beneficial to improving the accuracy of NN predictions. Inspired by Ref. [20], an iterative approach is introduced to find an optimal interval of the training set. By defining the initial training set as $[0,l^{(0)}]\cup[r^{(0)},1]$ with $l^{(0)}=0.1$ and $r^{(0)}=0.9$, the updating process can be started with the initial estimate critical probability $p_{c}^{(0)}$:
$l^{(i+1)}=\frac{l^{(i)}+p_{c}^{(i)}}{2},\quad r^{(i+1)}=\frac{r^{(i)}+p_{c}^{(i)}}{2},$ (7)
where $i$ represents the $i$-th expansion. $p_{c}^{(i)}$ is the estimate critical probability by SNN on the $i$-th expansion. For each iteration, it has to check whether at least 99% of the output similarity in the $i$-th interval of training is classified as category “0” or “1”. If the condition is met, it would continue the process, otherwise make a correction like $l^{(i+1)}\to l^{(i+1)}/2$. The iteration would be stopped when the interval of training set can’t be extended any more.
In Fig. 8, we show the iteration process to achieve the estimate critical value $p_{c}=0.6335$ of optimal training set $[0,0.58]\cup[0.69,1]$ for $L=32$, with anchor $p_{a}=0$. Obviously, the optimal training interval is much larger than the initial one, and the final result would also be more convinced. Table 2 gives the number of iterations, the optimal training interval and the critical probability predicted by SNN of $p_{a}=0$ for lattice size $L=8,16,32,48$ and $64$. Then, the critical probability of infinite size can also be obtained by the extrapolation in Fig. 9, as $p^{\infty}_{c}=0.6420$, and also by the data collapse we can get $\nu_{\perp}\simeq 1.09$. Table 3 shows the critical probabilities predicted by the optimal training interval under anchor point $p_{a}=0,0.4,0.8$ and $1$. Compared to the results of initial training interval in Table 1, a larger training set makes the SNN learn features more accurate, from which the results are less affected by the selected anchor values. Based on the previous summary, it can be inferred that by using the optimal training interval, selecting any anchor in training interval as test sample set can predict a reliable result. On the other hand, Table 3 also lists the results of supervised learning [47], autoencoder (AE) [47] and DANN [20] for comparison.
Figure 8: Evolution of the training set, and the corresponding critical bond probability values.
Table 2: The optimal training interval of $L=8,16,32,48$ and $64$, with iteration step and predicted $p_{c}$ at each size.
Lattice size Iteration step Optimal training interval $p_{c}$
L = 8 2 $[0,0.18]\cup[0.82,1]$ 0.6006
L = 16 6 $[0,0.53]\cup[0.73,1]$ 0.6258
L = 32 6 $[0,0.58]\cup[0.69,1]$ 0.6335
L = 48 6 $[0,0.59]\cup[0.67,1]$ 0.6323
L = 64 6 $[0,0.60]\cup[0.67,1]$ 0.6357
Figure 9: (a) Extrapolation of the critical probability $p_{c}$ to infinite lattice size with $p_{a}=0$, by the optimal training intervals in Table 3. (b) The collapse of the average output layer as a function of $(p-p_{c})L^{1/\nu_{\perp}}$.
Table 3: Comparing the supervised machine learning, autoencoder (AE), domain adversarial neural network (DANN), and SNN methods for calculating the critical value $p_{c}^{\infty}$ and spatial correlation exponent $\nu_{\perp}$ with optimal training intervals.
Method Supervised [47] AE [47] DANN [20] $p_{a}=0.0$ $p_{a}=0.4$ $p_{a}=0.8$ $p_{a}=1.0$
   $p_{c}^{\infty}$ 0.6408 0.643(2) 0.6453(5) 0.6420 0.6429 0.6452 0.6440
   $\nu_{\perp}$ 1.09(2) None 1.09(6) 1.09 1.08 1.1 1.09
5 Conclusion
In this paper, we have applied the Siamese Neural Network (SNN) on the (1+1)-dimensional bond directed percolation to study its critical behaviors of nonequilibrium system. Here the SNN is used as a semi-supervised learning method, in which the key information of critical behavior can be predicted by labeling only a part of raw configurations as input for training. And the output of the SNN is similarity, which transforms the classification problems of phase transition into a discussion of difference, broadening the application of Machine learning (ML) techniques in the study of critical behaviors.
By the output similarity of SNN, we predict the critical points of the system at different lattice sizes, and then use the fitting based on finite-size scaling to to obtain the theoretical values at unlimited system. Further, the curve of similarity also can be fit to a sigmoid-like function, which allows us to calculate the spatial correlation exponent $\nu_{\perp}$of non-equilibrium phase transition system through data collapse.
During the test, the selection of anchor points is of particular importance. By selecting the same training set, we discussed the impact of configurations under different anchors on the results. As the chosen anchor is outside the critical area, the predicted critical point increases with the value of anchor. For the anchor close to critical point, the curve of similarity shows a structure of single peak at the anchor, which indicates these configurations at the anchor only $100\%$ similar to themselves. In this case, the critical point cannot be predicted by the SNN. To make the calculation more efficient and reduce anchor dependence, we optimize the training interval through an iterative method. After optimization, the critical point result obtained by SNN is very accurate, comparable to the results of traditional supervised learning, AE and DANN.
With the widespread utilization of machine learning techniques in the fields of statistics and condensed matter physics, it is a promising future for further advancements. The SNN approach introduced in this research paper has the potential to offer novel insights into the exploration of DP and Ising-like phase transitions.
Acknowledgements
This work was supported in part by National Research Incubation Fund of Baoshan University(BYPY202216), the Baoshan University Doctoral Research Initiation Fund Project(BSKY202305), Academician workstation of Wenbangchun, Yunnan Province(Grant No.202205AF150032), Youth project of Yunnan Province Science and Technology Department(202401AU070035), Key Laboratory of Quark and Lepton Physics (MOE), Central China Normal University(Grant No.QLPL2022P01), and China Scholarship Council (CSC) - Swansea University scholarship.